Mickley Company’s plantwide predetermined overhead rate is $14.00 per direct labor-hour and its direct labor wage rate is $17.00 per hour. The following information pertains to Job A-500:

Direct materials $ 231
Direct labor $ 153

Required:

1. What is the total manufacturing cost assigned to Job A-500?

2. If Job A-500 consists of 40 units, what is the unit product cost for this job? (Round your answer to 2 decimal places.)
Answer #1
Concepts and reason

Cost: Cost is any value spent to produce a product or to render any service. The types of costs are fixed cost and variable cost.

Fixed cost: Fixed cost is a cost that remains same irrespective of the increase or decrease in the value of goods or any services rendered. It is the cost paid by the company that does not depend on the activities concerned with the business.

Variable cost: Variable cost is a cost that varies according to the output produced or any service rendered. It is the cost paid by the company that depends on the activities concerned with the business.

Job order costing: Job order costing involves appropriation of manufacturing or production cost to individual jobs. It is suitable for companies manufacturing heterogeneous products.

Fundamentals

Direct material cost: Direct material cost is the cost related to the purchase of the raw materials that are directly related to the production of the goods. It includes opening stock of materials, purchases, cost of purchases, and deducts the closing stock of materials.

Direct labor cost: It refers to the cost of providing wages to the workers who are directly associated with the production of goods or services rendered to the customers. The cost of direct labour includes the wages, payroll taxes, and all the benefits sponsored by the manufacturer.

Overhead cost: Overhead costs are the indirect cost incurred to operate the business. They are not connected with the production of goods or rendering any service. They do not include any direct cost.

Manufacturing overhead costs: The costs, which do not relate directly with the manufacturing of products, are referred to as manufacturing overhead costs or indirect costs.

Unit product cost: The cost ascertained and divided by the number of units of the product is the unit product cost. It includes prime cost, factory cost and production cost.

Direct professional labor hours: The labor hours spent for the actual production is known as direct professional labor hours. The activities of labor are directly involved with production.

1)

Determine the total manufacturing cost assigned to Job A-500.

Totalmanufacturingcost}=Directmaterials+Directlabor+(Laborhours×Predeterminedoverheadrate)=$231+$153+(9×$14)=$510\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Total manufacturing}}\\\\{\rm{cost}}\\\end{array} \right\}{\rm{ = Direct materials + Direct labor + }}\left( \begin{array}{l}\\{\rm{Labor hours}}\\\\ \times {\rm{Pre - determined overhead rate}}\\\end{array} \right)\\\\ = \$ 231 + \$ 153 + \left( {9 \times \$ 14} \right)\\\\ = \$ 510\\\end{array}

Therefore, the total manufacturing cost assigned to Job A-500 is $510.

Working notes:

Determine the labor hours:

Laborhours=DirectlaborcostLaborcostperunit=$153$17=9hours\begin{array}{c}\\{\rm{Labor hours = }}\frac{{{\rm{Direct labor cost}}}}{{{\rm{Labor cost per unit}}}}\\\\ = \frac{{\$ 153}}{{\$ 17}}\\\\ = 9{\rm{ hours}}\\\end{array}

Therefore, the labor hours are 9 hours.

2)

Determine the unit product cost:

Unitproductcost=TotalmanufacturingcostTotalunits=$51040=$12.75perunit\begin{array}{c}\\{\rm{Unit product cost = }}\frac{{{\rm{Total manufacturing cost}}}}{{{\rm{Total units}}}}\\\\ = \frac{{\$ 510}}{{40}}\\\\ = \$ 12.75{\rm{ per unit}}\\\end{array}

Therefore, the unit product cost for the job is $12.75 per unit.

Ans: Part 1

The total manufacturing cost assigned to Job A-500 is $510.

Part 2

The unit product cost for the job is $12.75 per unit.